1. if two dies are toced together what is the probability of having sum of 5 as the outcome 2.if two dies are toced together, what is the probability of having equal outcome 3.
Answer by ewatrrr(24785)

 
Hi,
36 outcomes
P(sum of five) = 4/36 = 1/9
P(If you meant pairs) = 6/36 = 1/6
___1_2_3__4__5__6
1|_2_3_4__5__6__7
2|_3_4_5__6__7__8
3|_4_5_6__7__8__9
4|_5_6_7__8__9_10
5|_6_7_8__9_10_11
6|_7_8_9_10_11_12